Catherine is a 35 year old mother of three young children. She was admitted to the hospital with multiple enterocutaneous fistulas 3 weeks ago, at which time she weighed 52 kg (116lbs). She is 165 cm (5ft 5in) tall. Catherine has a history of recurrent cervical cancer for which she had a hysterectomy 4 months before admission. During chemotherapy that followed, she had regular bouts with nausea and anorexia. Surgery was performed again. Her fistulas continued to drain for 2 weeks postoperatively, during which she tolerated clear liquids only. An intravenous drip of 10% glucose and 45% normal saline was ordered to supplement fluids and kcal. This week she developed peritonitis and has had a fever with a maximum temperature of 39 C (102 F) over the past 24 hours. Her weight has dropped to 41 kg (90lbs); drainage from the fistulas has become odorous. The patient was placed in isolation today and was advised by her physician that he intended to start her on total parenteral nutrition (TPN) and conduct some more tests to determine her progress.

Questions for Analysis:

Calculate Catherine's energy and protein needs, and account for increased needs.
Answer #1

Total energy need = BMR X activity factor

Here BMR for women = 655 + (9.6 X weight in kg) + (1.8 x height in cm) – (4.7 x age in yrs)

BMR = 655 + (9.6 x 41) + (1.8 x 165) — (4.7 x 35) = 655 + 393.6 + 297 — 164.5 = 1181.1calories/day

Activity factor for sedentary lifestyle with no exercise is 1.2

So total energy need = 1.2 x 1181.1 = 1417.32

So catherine requires 1417.32 calories per day

Protein need =

The lower limit for protein requirement = 0.8 x 41 = 32.8gm/day

The upper limit for protein requirement = 1.8 x 41 = 73.8gm/day

So we can say the protein requirement for Catherine ranges between 32.8 - 73.8gm/day.

The energy requirement has increased for Catherine because her body is recovering from a diseased conditions which requires for calorie and protein